OpenStudy (egbeach):

graph the equation 4. x = 2t – 1 y = t2 + 5, -4 ≤ t ≤ 4

OpenStudy (sdfgsdfgs):

x=2t - 1 so t=(x+1)/2 substitute t into the second eqn y=(x+1)^2/4 + 5 which represents a parabola. See the curve at wolframalpha: https://www.wolframalpha.com/input/?i=x+%3D+2t+%E2%80%93+1%2C+y+%3D+t%5E2+%2B+5